### v1

@@ 2.4 Vectorized fixed-point iteration
Following (9)–(10), we initialize the fixed-point iteration by approximating $Y_{n}^{j}$ for $j=1,\ldots,D$ as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j}.$ (13)
Then, $Y_{n}^{[k]}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
$\begin{split}F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k-1]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=1,\ldots,D,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (14)
@@
If $Y_{n}^{j,[k]}=Y_{n}^{j,[k-1]}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $Y_{n}^{j,[k+1]}=Y_{n}^{j,[k]}$ would hold for all $j\in\{1,\ldots,D\}$. However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define $\Delta_{n}^{j,[k]}=\|Y_{n}^{j,[k]}-Y_{n}^{j,[k-1]}\|_{\infty}$. The iteration will stop after the $k$th iteration if the following condition is met for all $j\in\{1,\ldots,D\}$:
@@
Systems of second Order Differential Equations
We now consider systems of second-order differential equations of the form
$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),\quad q(t_{0})=q_{0},\quad v(t_{0})=v_{0},$ (16)
where $q$ and $v$ are vectors with $d$ components each. Hamiltonian systems with a Hamiltonian function of the form
$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (17)
can be expressed in this form via the linear change of variables $p=Mv$.
System (16) is a specific instance of the general form (1), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$. It is known [18, 13], that a partitioned form of fixed-point iteration can generally reduce the number of iterations needed for convergence. Such partitioned fixed-point iteration also admits a vectorized implementation that we now describe.
We initialize the fixed-point iteration as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j},\quad j=d+1,\ldots,2d.$ (18)
Then, for $k=1,2,\ldots$,
$\begin{split}L_{n}^{j,[k]}&=h\,(b\circ Y_{n}^{d+j,[k-1]}),\quad j=1,\ldots,d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,d,\\ F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=d+1,\ldots,2d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=d+1,\ldots,2d.\end{split}$ (19)
The process will stop after the $k$th iteration if (15) holds for all $j\in\{1,\ldots,d\}$.
3 Implementation aspects of IRKGL16-SIMD
@@ 3.1 SIMD-paralellization
3.2 Julia language and SIMD.jl
@@
Our SIMD-vectorized implementation of IRKGL16 (IRKGL16-SIMD) is written in Julia language and relies on the SIMD.jl package. Julia is a high-level dynamic language that allows programmers to write clear, high-level, generic and abstract code resembling mathematical formulas, yet produces fast, low-level machine code that has traditionally only been generated by static languages [3, 4]. By default, Julia employs just-in-time (JIT) compilation, generating LLVM intermediate code that the LLVM compiler framework then translates into optimized machine code.
@@
In Julia, there are several ways to explicitly apply SIMD vectorization, and the SIMD.jl package [19] provides a convenient solution for this purpose. The SIMD.jl library introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T. In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16-SIMD), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type ‘Float64‘ (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}. The standard arithmetic and logical operations are designed to be applied element-wise in parallel, producing a SIMD vector as the result. SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized SIMD instructions that utilize the SIMD registers available on the target platform.
3.3 VecArray and IRKGL16-SIMD
@@
The vectorized form of fixed-point iteration (13)–(14) (resp., (18)–(19)) for first-order problems (1) (resp., second-order problems (16)) is well-suited for SIMD implementation. SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (13)–(14) and (18)–(19) can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
@@
In addition, the evaluation of $F(t_{n-1}+ch,Y_{n}^{[k-1]})$ in (14) and (19) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16-SIMD, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl. Such seamless SIMD parallelization is handled automatically with the help of SIMD.jl and an additional parametrized abstract type, VecArray{s,T}, that we have implemented to efficiently represent and handle abstract arrays of elements of type Vec{s,T} (where T is either Float64 or Float32). The $s$-vectors $b,c,\mu_{i},\nu_{i}$, $Y_{n}^{j}$, $F_{n}^{j}$, and $L_{n}^{j}$ considered in Subsection 2.4 are represented in IRKGL16-SIMD as vectors of type Vec{s,T}, while $Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})$, $F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})$, and $L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})$, and also $\mu=(\mu_{1},\ldots,\mu_{s})$ and $\nu=(\nu_{1},\ldots,\nu_{s})$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
Let us consider for instance the VecArray object Yn representing $Y_{n}$:
@@
  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $Y_{n}^{j}$.
@@
  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $Y_{n}^{j}$. Here, we are assuming that V is a vector of type Vec{s,T}.
@@
Listing LABEL:code:ex1 shows the section of the IRKGL16-SIMD code where a single vectorized fixed-point iteration (14) is implemented.
@@ 10 Yn[j]=yn[j]+dYnj
The generic function f_ODE!(dy, y, parms, t) in Listing LABEL:code:ex2 supports a wide range of argument types. For example, it can be called with t as a scalar of type T = Float64 (or alternatively T=Float32 or T=BigFloat), and dy and y as standard Julia arrays containing elements of type T. Additionally, the function accepts other argument types, provided that the operations within the f_ODE! definition are valid for them. Specifically, the function call f_ODE!(Fn, Yn, parms, tn + h*c) in Listing LABEL:code:ex1, where Fn and Yn are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and tn + h*c is of type Vec{s, T}, is equivalent to the first line of (14), representing the parallel evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$.
This design enables parallel evaluations of the ODE system’s right-hand side in a way that is transparent to the user, provided that the right-hand side $f(t,y)$ of the ODE system is defined as a generic function f_ODE!(dy, y, parms, t) without branching in terms of arithmetic operations and elementary functions compatible with SIMD.jl. This requirement is met for all the problems considered in the numerical experiments in Section 5 below.
However, not all ODE systems (1) support a generic implementation of f_ODE!(dy, y, parms, t) that enables seamless $s$-fold parallel evaluations of the system’s right-hand side. For instance, this may not be feasible if the definition of $f(t,y)$ involves unavoidable branching, which disrupts the uniformity needed for SIMD processing. In such cases, however, IRKGL16-SIMD can still be applied if the user provides a specialized implementation of f_ODE!(dy, y, parms, t) for instances where dy and y are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and t is of type Vec{s, T}.
@@
In some cases, such as the $N$-body example considered in Section 5, it is more convenient for the user to organize the state variables in a higher-dimensional array rather than a one-dimensional array. To accommodate this, we have parameterized our abstract array type with an additional parameter dim, as VecArray{s, T, dim}. For one-dimensional arrays of vectors of type Vec{s, T}, we set dim = 2, with the underlying data stored in a two-dimensional standard array of type Array{T, 2}. More generally, VecArray{s, T, dim} refers to $(\texttt{dim}-1)$-dimensional abstract arrays of vectors of type Vec{s, T}, with the data stored in a dim-dimensional array of type Array{T, dim}. Our implementations of the getindex and setindex! functions work as expected in this more general case, allowing users to apply IRKGL16-SIMD with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
@@ 4 State-of-the-art explicit symplectic integrators
In the context of Hamiltonian systems that can be decomposed into two (or more) exactly solvable components, explicit symplectic integrators can be constructed via operator splitting techniques. A widely used approach is the second-order Strang splitting method [21], which yields a symmetric and symplectic scheme by composing the flows associated with the individual sub-Hamiltonians. Specifically, the Strang splitting scheme advances the solution over a time step $h$ via the composition
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h}\circ\varphi^{A}_{h/2}.$ (21)
This method is second-order accurate, time-reversible, and symplectic. The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the exact flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
For instance, consider a separable Hamiltonian of the form (17), where $H_{A}(q,p)=\tfrac{1}{2}p^{T}M^{-1}p$ and $H_{B}(q,p)=U(q)$. In this case, the exact flows are given by $\varphi^{A}_{t}(q,p)=(q+tM^{-1}p,\,p)$ and $\varphi^{B}_{t}(q,p)=(q,\,p-t\nabla U(q))$.
@@
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h/2}\circ\varphi^{C}_{h}\circ\varphi^{B}_{h/2}\circ\varphi^{A}_{h/2}.$ (22)
@@
$\psi_{h}=\phi_{\gamma_{s}h}\circ\phi_{\gamma_{s-1}h}\circ\dots\circ\phi_{\gamma_{2}h}\circ\phi_{\gamma_{1}h}$ (23)
@@
We consider the following symmetric compositions of time-symmetric second-order schemes of order $r=2,4,6,8,10$ recommended in [7]:
@@
    Strang splitting: $r=2$th order with $s=1$ stage.
@@
    SUZ90: $r=4$th order with $s=5$ stages proposed by Suzuki(1990) [22].
@@
    SS05: a sixth-order scheme with $s=13$ stages, a eighth-order one with $s=21$ and a tenth-order one with $s=35$ proposed by Sofroniu and Spalleta(2005) [20].
@@
$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}\circ\varphi_{a_{s}h}^{A}\circ\dots\circ\varphi_{b_{1}h}^{B}\circ\varphi_{a_{1}h}^{[A]}$ (24)
@@
Clearly, (24) cannot be written in the form (23) for arbitrary sequences of coefficients $a_{j},b_{j}$. Such more general family of methods are of particular interest in the case of Hamiltonians of the form (17), for which efficient high-order schemes have been constructed. They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (16[18]. We consider the following splitting integrators of RKN type of order $r=6$ and $r=8$ recommended in [7]:

### v2

@@ 2.5 Stage-wise vectorization of fixed-point iteration
Following (6)–(7), we initialize the fixed-point iteration by approximating $\boldsymbol{Y_{n}^{j}}$ for $j=1,\ldots,D$ as
$\boldsymbol{Y_{n}^{j,[0]}}=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\nu_{i}}\,L_{n-1,i}^{j}.$ (24)
Then, $\boldsymbol{Y_{n}^{[k]}}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
$\begin{split}\boldsymbol{F_{n}^{[k]}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},\boldsymbol{Y_{n}^{[k-1]}}),\\ \boldsymbol{L_{n}^{j,[k]}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j,[k]}}),\quad j=1,\ldots,D,\\ \boldsymbol{Y_{n}^{j,[k]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}\,L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (25)
This formulation naturally enables SIMD implementation: the $s$-vectors can be stored in SIMD registers, and operations like $\boldsymbol{b}*\boldsymbol{F_{n}^{j}}$ correspond to single SIMD instructions.
As for the alternative IRKN form (19)–(21) of IRKGL methods for second order ODE systems, we initialize the fixed-point iteration as
$\begin{split}\boldsymbol{Y_{n}^{j,[0]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\nu_{i}}\,L_{n-1,i}^{j},\quad j=d+1,\ldots,2d,\\ \boldsymbol{L_{n}^{j,[0]}}&=h\,(\boldsymbol{b}*\boldsymbol{Y_{n}^{d+j,[0]}}),\quad j=1,\ldots,d,\\ \boldsymbol{Y_{n}^{j,[0]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}L_{n,i}^{j,[0]},\quad j=1,\ldots,d.\end{split}$ (26)
Then, for $k=1,2,\ldots$,
$\begin{split}\boldsymbol{F_{n}^{[k]}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},Y_{n}^{[k-1]}),\\ \boldsymbol{L_{n}^{j,[k]}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j,[k]}}),\quad j=d+1,\ldots,2d,\\ \boldsymbol{Y_{n}^{j,[k]}}&=y_{n-1}^{j}+h\,\left(\boldsymbol{c}\,y_{n-1}^{d+j}+\sum_{i=1}^{s}\boldsymbol{\eta_{i}}\,L_{n,i}^{d+j,[k]}\right),\quad j=1,\ldots,d,\end{split}$ (27)
@@
If $\boldsymbol{Y_{n}^{j,[k]}}=\boldsymbol{Y_{n}^{j,[k-1]}}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $\boldsymbol{Y_{n}^{j,[k+1]}}=\boldsymbol{Y_{n}^{j,[k]}}$ would hold for all $j\in\{1,\ldots,D\}$. However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define
$\Delta_{n}^{j,[k]}=\|\boldsymbol{Y_{n}^{j,[k]}}-\boldsymbol{Y_{n}^{j,[k-1]}}\|_{\infty}.$ (28)
The iteration will stop after the $k$th iteration if the following condition is met for all $j\in\{1,\ldots,D\}$:
@@
For second order ODE systems, condition (29) is checked for all $j\in\{1,\ldots,d\}$.
3 Implementation aspects of IRKGL16
@@ 3.1 SIMD-paralellization
3.2 Julia language and the package SIMD.jl
@@
Our SIMD-vectorized implementation of IRKGL16 is written in Julia language and relies on the package SIMD.jl [22]. Julia is a high-level dynamic language that allows programmers to write clear, high-level, generic and abstract code resembling mathematical formulas, yet produces fast, low-level machine code that has traditionally only been generated by static languages [4, 5]. By default, Julia employs just-in-time (JIT) compilation, generating LLVM intermediate code that the LLVM compiler framework then translates into optimized machine code.
@@
In Julia, there are several ways to explicitly apply SIMD vectorization, and the package SIMD.jl provides a convenient solution for this purpose. That package introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T. In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type Float64 (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}. The standard arithmetic and logical operations are designed to be applied element-wise in parallel, producing a SIMD vector as the result. The package SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized instructions that utilize the SIMD registers available on the target platform, while also enabling other low-level optimizations (e.g., instruction-level parallelism) afforded by the Vec{s,T} abstraction.
3.3 VecArray and IRKGL16
@@
The vectorized form of fixed-point iteration (24)–(25) (resp., (26)–(27)) for first-order problems (2) (resp., second-order problems (10)) is well-suited for SIMD implementation. SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (24)–(25) and (26)–(27), as well as the computations in (28), can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
@@
In addition, the evaluation of $F(t_{n-1}+\boldsymbol{c}\,h,\boldsymbol{Y_{n}^{[k-1]}})$ in (25) and (27) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl. Such seamless SIMD parallelization is handled automatically with the help of SIMD.jl and an additional parametrized abstract type, VecArray{s,T}, that we have implemented to efficiently represent and handle abstract arrays of elements of type Vec{s,T} (where T is either Float64 or Float32). The $s$-vectors $\boldsymbol{b,c,\mu_{i},\eta_{i},\nu_{i}}$, $\boldsymbol{Y_{n}^{j}}$, $\boldsymbol{F_{n}^{j}}$, and $\boldsymbol{L_{n}^{j}}$ considered in Subsection 2.5 are represented in IRKGL16 as vectors of type Vec{s,T}, while $\boldsymbol{Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})}$, $\boldsymbol{F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})}$, and $\boldsymbol{L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})}$, and also $\boldsymbol{\mu=(\mu_{1},\ldots,\mu_{s})}$, $\boldsymbol{\eta=(\eta_{1},\ldots,\eta_{s})}$, and $\boldsymbol{\nu=(\nu_{1},\ldots,\nu_{s})}$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
Let us consider for instance the VecArray object Yn representing $\boldsymbol{Y_{n}}$:
@@
  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $\boldsymbol{Y_{n}^{j}}$.
@@
  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $\boldsymbol{Y_{n}^{j}}$. Here, we are assuming that V is a vector of type Vec{s,T}.
@@
Listing LABEL:code:ex1 shows the section of the IRKGL16 code where a single vectorized fixed-point iteration (25) is implemented.
@@ 10 Yn[j]=yn[j]+dYnj
The generic function f_ODE!(dy, y, parms, t) in Listing LABEL:code:ex2 supports a wide range of argument types. For example, it can be called with t as a scalar of type T = Float64 (or alternatively T=Float32 or T=BigFloat), and dy and y as standard Julia arrays containing elements of type T. Additionally, the function accepts other argument types, provided that the operations within the f_ODE! definition are valid for them. Specifically, the function call f_ODE!(Fn, Yn, parms, tn + h*c) in Listing LABEL:code:ex1, where Fn and Yn are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and tn + h*c is of type Vec{s, T}, is equivalent to the first line of (25), representing the parallel evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$. This design enables parallel evaluations of the ODE system’s right-hand side in a manner that is transparent to the user, provided that the right-hand side $f(t,y)$ is implemented as a generic function f_ODE!(dy, y, parms, t) without branching and using only arithmetic operations and elementary functions compatible with SIMD.jl. All test problems considered in Section 5 satisfy this requirement.
However, not all ODE systems (2) admit a generic implementation of f_ODE!(dy, y, parms, t) that supports seamless $s$-fold parallel evaluation of the system’s right-hand side. In particular, this may fail when $f(t,y)$ contains unavoidable branching, which breaks the uniform instruction flow required for SIMD processing. In such cases, the right-hand side must be evaluated sequentially for each stage vector $Y_{i}$, although the remaining vectorized computations in (24)–(25) (for first-order ODEs), (26)–(27) (for second-order ODEs) and (28) remain applicable.
In the current implementation of IRKGL16, simultaneous SIMD evaluation of the user-supplied function $f$ is performed only when the integrator option fseq=false is explicitly selected by the user. All other computations are executed using SIMD vectorization whenever the state variables are of type Float32 (single precision) or Float64 (double precision). A fully sequential fallback implementation is used for state variables of other types (e.g., BigFloat) or whenever the simd option is explicitly disabled (simd=false) by the user.
@@
In some cases, such as the $N$-body example considered in Section 5, it is more convenient for the user to organize the state variables in a higher-dimensional array rather than a one-dimensional array. To accommodate this, we have parameterized our abstract array type with an additional parameter dim, as VecArray{s, T, dim}. For one-dimensional arrays of vectors of type Vec{s, T}, we set dim = 2, with the underlying data stored in a two-dimensional standard array of type Array{T, 2}. More generally, VecArray{s, T, dim} refers to $(\texttt{dim}-1)$-dimensional abstract arrays of vectors of type Vec{s, T}, with the data stored in a dim-dimensional array of type Array{T, dim}. Our implementations of the getindex and setindex! functions work as expected in this more general case, allowing users to apply IRKGL16 with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
@@ 4 State-of-the-art explicit symplectic integrators
In the context of Hamiltonian systems that can be decomposed into two (or more) exactly solvable components, explicit symplectic integrators can be constructed via operator splitting techniques. A widely used approach is the second-order Strang splitting method [24], which yields a symmetric and symplectic scheme by composing the $t$-flow maps $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ associated with the individual sub-Hamiltonians. Specifically, the Strang splitting scheme advances the solution over a time step $h$ via the composition
$\phi_{h}=\varphi^{A}_{h/2}*\varphi^{B}_{h}*\varphi^{A}_{h/2}.$ (31)
This method is second-order accurate, time-reversible, and symplectic. The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the $t$-flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
For instance, consider a separable Hamiltonian of the form (8), where $H_{A}(q,p)=\tfrac{1}{2}p^{T}M^{-1}p$ and $H_{B}(q,p)=U(q)$. Via the linear change of variables $p=Mv$, the exact flows are given by
$\varphi^{A}_{t}(q,v)=(q+t\,v),\quad\varphi^{B}_{t}(q,v)=(q,\,v-t\,g(q)),$ (32)
where $g(q)=M^{-1}\nabla U(q)$.
@@
$\phi_{h}=\varphi^{A}_{h/2}*\varphi^{B}_{h/2}*\varphi^{C}_{h}*\varphi^{B}_{h/2}*\varphi^{A}_{h/2}.$ (33)
@@
$\psi_{h}=\phi_{\gamma_{s}h}*\phi_{\gamma_{s-1}h}*\dots*\phi_{\gamma_{2}h}*\phi_{\gamma_{1}h}$ (34)
@@
We consider the following symmetric compositions of time-symmetric second-order schemes of orders $r=2,4,6,8,10$, as recommended in [8]:
@@
    Strang splitting: a second-order method ($r=2$) with $s=1$ stage.
@@
    SUZ90: a fourth-order method ($r=4$) with $s=5$ stages, proposed by Suzuki (1990) [26].
@@
    SS05: a sixth-order scheme with $s=13$ stages, an eighth-order one with $s=21$ stages, and a tenth-order one with $s=35$ stages proposed by Sofroniou and Spalleta (2005) [23].
@@
$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}*\varphi_{a_{s}h}^{A}*\dots*\varphi_{b_{1}h}^{B}*\varphi_{a_{1}h}^{[A]}$ (35)
@@
Clearly, (35) cannot be written in the form (34) for arbitrary sequences of coefficients $a_{j},b_{j}$. Such more general family of methods are of particular interest in the case of Hamiltonians of the form (8), for which efficient high-order schemes have been constructed. They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (10[21]. We consider the following splitting integrators of RKN type of order $r=6$ and $r=8$ recommended in [8]: